OpenStudy (cutiecomittee123):

what is the minimum value of the function y=-1+6cos(2pi/7(x-5))

OpenStudy (anonymous):

just graph the function..

OpenStudy (cutiecomittee123):

I dont have my calculator and I dont want to graph it, is there any other way to calculate it?

OpenStudy (anonymous):

yes

OpenStudy (baru):

the lowest value that 'cos' can take is -1. so the the lowest value the function can take is y=-1 + 6*(-1) =-7

OpenStudy (anonymous):

not that \[-1\le \cos(Z) \le1 \]

OpenStudy (anonymous):

so\[-1 \le \cos(\frac{ 2 \pi }{ 7 }(x-5))\le1\] \[-6 \le 6\cos(\frac{ 2 \pi }{ 7 }(x-5))\le6\] \[-7 \le -1+6\cos(\frac{ 2 \pi }{ 7 }(x-5))\le7\]

OpenStudy (anonymous):

this tells us that our minimum value for this function is -7 and our max value os 5

OpenStudy (anonymous):

the last line should be \[-7 \le -1+6\cos(\frac{ 2 \pi }{ 7 }(x-5))\le5\]

OpenStudy (anonymous):

something along those lines..

OpenStudy (ybarrap):

$$ y=-1+6\cos(2\pi/7(x-5)) $$ Break it down The smallest that \(6\cos(2\pi/7(x-5))\) can be is -6. See that? -1 is a constant and the lowest it can be is -1. So, the lowest y will be is -1 + (-6) = -7. Does this make sense?